A horse pulled 500N to a distance of 12 m in 3 sec at a constant speed. What was the power of the horse in

a ) watts b) horsepower c) ft - lb d) kilowatts

s


1
Expert's answer
2021-05-07T13:54:09-0400

(a)

"P=\dfrac{W}{t}=\dfrac{Fd}{t},""P=\dfrac{500\ N\cdot12\ m}{3\ s}=2000\ W."

(b)

"P=2000\ W\cdot\dfrac{1\ hp}{746\ W}=2.68\ hp."

(c)

"P=\dfrac{500\ N\cdot\dfrac{0.2248\ lb}{1\ N}\cdot12\ m\cdot\dfrac{3.281\ ft}{1\ m}}{3\ s}=1475\ \dfrac{ft\cdot lb}{s}."

(d)

"P=2000\ W \cdot\dfrac{1\ kW}{1000\ W}=2\ kW."